The average bill for car repairs at a car service center is $196, with a standard deviation of $44. Assuming the bills to be normally distributed, find the probability of a bill exceeding $300.
Answer #1
for normal distribution z score =(X-μ)/σ
here mean=       μ= 196
std deviation   =σ= 44.0000

probability of a bill exceeding $300 =P(X>300)=P(Z>(300-196)/44)=P(Z>2.36)=0.0091